◯^{R}×

?? 1/pci104[width =

55mm]anexos/anexo1/imagens_{1}/pci1041[P C−

104]P C − 104V istadecimaevistadebaixorespectivamente. re- set jumper

?? FireWire IEEE 1394 jumpers 1/pci104eP lacaexpansodoP C − 104 Unetbootin

sudo aptget install unetbootin

?? 1/unetbootin[Unetbootin]Unetbootin.

Lightweight X11 Desktop

En- vi- ron- ment ?? 1/unetbootin1[Selecodeopes]Selecodeopes.

?? 1/unetbootin3[F imdacriaodaimagemUSB]F imdacriaodaimagemUSB.

?? 1/lubuntu1[Menuselecoidioma]Menuselecoidioma.

1/lubuntu2Menuconfiguraoinstalao.

?? ?? 1/lubuntu3[Menuconfiguraodisco]Menuconfiguraodisco.

?? 1/lubuntu4[Menuselecofusohorrio]Menuselecofusohorrio.

?? 1/lubuntu5[Menuselecodisposioteclado]Menuselecodisposioteclado.

?? 1/lubuntu6[Menuconfiguraoutilizador]Menuconfiguraoutilizador.

?? 1/lubuntu7[Instalaosistemaoperativo]Instalaosistemaoperativo.

?? 1/lubuntu8[AmbientegrficoLXDE]AmbientegrficoLXDE. swap

software Solid State Drive Swap ? Swap GParted GParted

sudo aptget install gparted

GParted Swap updatemanager updatemanager

?? 1/upman[GestordeActualizaes]GestordeActualizaes.

Firefly Fly-Cap-ture 2.0 www.ptgrey.com

tar - xvf flycapture2-2.1.3.4-i386-pkg.tgz

sudo aptget install libraw1394-11 libgtk2.0-0 libgtkmm-2.4-dev

sudo aptget install libglademm-2.4-dev